# ILID: Native Script Language Identification for Indian Languages

### v1

@@ Appendix B ISO 639-2 Codes
Appendix C Results for Individual Languages
The language wise performance of the models are shown in Tables 6 and 7.
Language LR DT RF SVM SGD k-NN NB ADABoost MURIL FastText
Assamese (asm) 0.99 0.97 0.99 0.99 0.98 0.91 0.98 0.77 0.99 0.99
Bengali (ben) 1.00 0.99 0.99 1.00 0.99 0.93 0.99 0.92 1.00 1.00
Bodo (brx) 0.96 0.92 0.96 0.97 0.94 0.90 0.96 0.68 0.97 0.96
Dogri (doi) 0.90 0.82 0.89 0.90 0.88 0.71 0.90 0.19 0.95 0.92
English (eng) 1.00 0.99 1.00 1.00 1.00 1.00 0.99 0.95 1.00 1.00
Konkani (gom) 0.97 0.85 0.95 0.97 0.93 0.80 0.96 0.29 0.98 0.98
Gujarati (guj) 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.97 1.00 1.00
Hindi (hin) 0.99 0.93 0.99 0.99 0.98 0.83 0.94 0.79 1.00 0.99
Kannada (kan) 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.96 1.00 1.00
Kashmiri (kas) 0.99 0.95 0.97 0.99 0.94 0.97 0.98 0.75 1.00 1.00
Maithili (mai) 0.87 0.78 0.84 0.86 0.84 0.67 0.84 0.26 0.92 0.90
Malayalam (mal) 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.96 1.00 1.00
Marathi (mar) 0.98 0.84 0.96 0.98 0.93 0.76 0.94 0.52 0.99 0.99
Manipuri (mni) 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.90 0.72 0.72
Nepali (npi) 0.92 0.84 0.90 0.91 0.86 0.76 0.89 0.28 0.94 0.92
Odia (ory) 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.90 1.00 1.00
Punjabi (pan) 1.00 0.99 1.00 1.00 0.99 0.99 0.99 0.91 1.00 1.00
Sanskrit (san) 0.97 0.92 0.99 0.98 0.97 0.84 0.94 0.78 0.99 0.99
Santali (sat) 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.95 0.71 0.71
Sindhi (snd) 0.94 0.86 0.92 0.93 0.90 0.80 0.94 0.24 0.96 0.94
Tamil (tam) 1.00 0.99 1.00 1.00 0.99 0.99 0.99 0.95 1.00 1.00
Telugu (tel) 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.91 1.00 1.00
Urdu (urd) 0.99 0.94 0.96 0.99 0.96 0.97 0.98 0.76 1.00 1.00
Table 6: Dev dataset report: word+char combined Classifier performance (F1-scores) across 22 languages. Bold indicates top-performing model(s) per language.
Language LR DT RF SVM SGD k-NN NB AdaBoost MuRIL FastText
Assamese (asm) 0.99 0.96 0.98 0.99 0.98 0.91 0.97 0.78 0.99 0.99
Bengali (ben) 1.00 0.98 0.99 1.00 0.99 0.93 0.99 0.93 1.00 1.00
Bodo (brx) 0.95 0.90 0.96 0.96 0.95 0.89 0.96 0.69 0.97 0.96
Dogri (doi) 0.91 0.82 0.91 0.90 0.88 0.71 0.91 0.21 0.94 0.92
English (eng) 1.00 1.00 0.99 1.00 0.99 1.00 0.99 0.96 1.00 1.00
Konkani (gom) 0.97 0.86 0.94 0.97 0.92 0.81 0.95 0.31 0.98 0.97
Gujarati (guj) 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.97 1.00 1.00
Hindi (hin) 0.99 0.93 0.98 0.99 0.97 0.84 0.94 0.74 0.99 0.99
Kannada (kan) 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.96 1.00 1.00
Kashmiri (kas) 0.98 0.93 0.97 0.98 0.94 0.98 0.98 0.74 1.00 0.99
Maithili (mai) 0.85 0.78 0.84 0.85 0.81 0.67 0.82 0.26 0.91 0.89
Malayalam (mal) 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.97 1.00 1.00
Marathi (mar) 0.97 0.85 0.95 0.98 0.92 0.79 0.94 0.50 0.98 0.98
Manipuri (mni) 1.00 1.00 1.00 1.00 1.00 0.99 1.00 0.89 0.73 0.72
Nepali (npi) 0.90 0.84 0.90 0.90 0.86 0.76 0.89 0.28 0.94 0.91
Odia (ory) 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.90 1.00 1.00
Punjabi (pan) 1.00 0.99 0.99 1.00 0.99 0.99 1.00 0.91 1.00 1.00
Sanskrit (san) 0.98 0.92 0.99 0.98 0.97 0.84 0.94 0.79 0.98 0.98
Santali (sat) 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.94 0.72 0.71
Sindhi (snd) 0.94 0.87 0.93 0.93 0.90 0.80 0.92 0.23 0.96 0.94
Tamil (tam) 1.00 0.99 1.00 1.00 0.99 1.00 1.00 0.96 1.00 1.00
Telugu (tel) 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.92 1.00 1.00
Urdu (urd) 0.98 0.92 0.97 0.99 0.96 0.98 0.98 0.75 1.00 0.99
Table 7: Test dataset report showing word+char combined classifier performance (F1-scores) across 22 languages. Bold indicates top-performing model(s) per language.

### v2

@@ 3.3 Sampling From Existing Corpora
lang #sents #words #chars avg_word_len avg_sent_len #unique_words TTR_words
mai 10000 65884 288223 3.526 28.822 6743 0.102
eng 10000 225036 1052283 3.721 105.228 22709 0.101
kan 10000 122400 861095 6.117 86.109 34644 0.283
san 10000 130353 858829 5.665 85.883 50506 0.387
gom 10000 58282 296807 4.264 29.681 8104 0.139
hin 10000 106402 476919 3.576 47.692 8451 0.079
mar 10000 89232 526021 5.007 52.602 19470 0.218
asm 10000 64251 311313 4.001 31.131 7196 0.112
urd 10000 97233 387031 3.083 38.703 19987 0.206
tel 10000 81281 524216 5.572 52.422 19554 0.241
tam 10000 144307 1018268 6.126 101.827 45912 0.318
snd_Arab 10000 113249 479446 3.322 47.945 7409 0.065
ben 10000 170048 877189 4.217 87.719 30007 0.176
snd_Deva 10000 70743 312570 3.56 31.257 6249 0.088
ory 10000 91919 526927 4.841 52.693 5400 0.059
npi 10000 59095 296473 4.186 29.647 7966 0.135
mni_Mtei 10000 73043 318189 3.493 31.819 7245 0.099
mal 10000 117447 985898 7.48 98.59 37010 0.315
doi 10000 75469 313613 3.288 31.361 5904 0.078
pan 10000 168450 721285 3.341 72.129 18032 0.107
brx 10000 65717 343731 4.383 34.373 7828 0.119
sat 10000 78554 360814 3.72 36.081 4067 0.052
mni_Beng 10000 186109 1208020 5.545 120.802 27350 0.147
guj 10000 175966 911173 4.235 91.117 28195 0.16
kas 10000 69270 297326 3.437 29.733 7752 0.112
Table 4: Language Wise Corpus Statistics Where the column TTR_words refers to the Type Token Ratio for Words
Model KNN DT RF SVM NB LogReg SGD ILID Bh-Ab
Dev Test
Voting-1 1.00 0.96 1.00
Voting-2 1.00 0.96 1.00
Voting-3 1.00 0.96 1.00
Voting-4 0.99 0.96 1.00
Voting-5 0.99 0.99 1.00
MuRIL 0.96 0.96 0.9
FastText 0.96 0.96 0.9
Table 5: Comparison of Voting Classifiers, MuRIL, and FastText on ILID dataset and Bhasha-Abhijnaanam (Bh-Ab) dataset. Highest scores are marked in bold.
3.4 Corpora Statistics
After collecting and cleaning the dataset, we compute several corpora statistics. These statistics are excellent indicators of the levels and complexities of each language. The properties of the corpora considered include the total number of words, the total number of characters, the total number of unique words, average sentence length, the average word length, and the type / token ratio textinspector (2025) which are markers of the representativeness of the corpora. These statistics are shown in Table 4.
4 ILID Model
The ILID model is a classifier that can categorize a piece of text into one of the 25 classes that includes English and 22 Indian languages. The architecture of the proposed system is shown in Figure 1. We implement three types of approach for designing the classifiers.
4.1 Machine Learning Models
In this approach, each sentence is represented by a TF-IDF Sparck Jones (1972) vector in a bag-of-words setting. We utilize both word-level and character-level TF-IDF representations to enhance textual feature extraction.
4.1.1 Word-Level TF-IDF Based Classification
[…]
4.1.2 Character-Level TF-IDF Based Classification
To overcome the limitations of word-level models, we introduce character-level TF-IDF representations using character n-grams ranging from 2 to 6. This approach captures sub-word structures, prefixes, suffixes, and root forms that are particularly effective in handling typographical errors, morphological variations, and out-of-vocabulary words. Character-level features are especially helpful in language identification tasks involving noisy or informal text, enabling the model to learn finer linguistic patterns.
4.1.3 Combined Feature-Based Classification
To leverage the strengths of both word and character level features, we develop a combined TF-IDF representation by concatenating the two feature sets into a single vector. This hybrid approach provides a more robust and language-agnostic representation. We train eight traditional machine learning classifiers—Support Vector Machine (SVM), Logistic Regression (LR), Random Forest (RF), Decision Tree (DT), Naive Bayes (NB), Stochastic Gradient Descent (SGD), and K-Nearest Neighbors (KNN)—on these features. Furthermore, we explore various ensemble models involving combinations of 3, 4, and 5 diverse base classifiers from the pool of classifiers. Due to space constraints, only the top five performing ensembles are selected based on F1 scores on the development and test sets, are presented in Table 5.
4.2 FastText Classifier
FastText Joulin et al. (2016b, a) utilizes word embeddings Bojanowski et al. (2017) composed of character n-grams that better represent rare words and orthographic similarities. FastText classifier is a linear classifier and is very fast in computation. It can generate embeddings for out-of-vocabulary (OOV) words, which are common in user-generated or noisy text data.
4.3 Pretrained BERT Model
Pretrained subword based contextual language models such as BERT Devlin et al. (2019), XLM Conneau et al. (2020), RoBERTa Zhuang et al. (2021) have proven to be very effective in text classification tasks or generally natural language understanding tasks. One variant of BERT, MuRIL Khanuja et al. (2021) is pretrained on large amounts of corpora in Indian languages. The pre-training data is also augmented with translated and transliterated texts that makes the model capture cross-lingual and code-mixed embeddings more efficiently. Hence, we fine-tune the MuRIL pretrained model on the ILID train set for our task.
5 Experimental Details
The machine learning models have been implemented using the Scikit-learn Pedregosa et al. (2011) framework. Similarly, the FastText library  [^5] https://fasttext.cc/docs/en/supervised-tutorial.html is utilized to implement the language identifiers of the texts modeled as text classifiers Joulin et al. (2016b, a). The MuRIL Khanuja et al. (2021) models are fine-tuned using the Huggingface transformer Wolf et al. (2019) framework on an NVIDIA H100 GPU with 94GB RAM. The batch size, the number of epochs, the learning rate, maximum sequence length, and weight decay are set to 32, 10, 0.00002, 256, and 0.01 respectively. The fine-tuned MuRIL model is available at https://huggingface.co/pruthwik/ilid-muril-model.
6 Evaluation Metrics
All models have been evaluated using the macro F1 scores. The macro F1 score averages the F-1 scores across all languages. In order to evaluate the score the F1-score of each language, the precision and recall scores are also computed at the macro level. Each ensemble employs a voting mechanism to determine the predicted labels. Two types of voting mechanism are used: soft and hard. Hard voting is based on majority voting among the classifiers present in the ensemble. This type of voting is performed when the SVM classifier is part of the ensemble because SVM is inherently a non-probabilistic classifier. Soft voting aggregates the probability estimates of individual classifiers in the ensemble and the class with argmax of the sums of the predicted probabilities is chosen as the label. Ensembles where SVM is not included are evaluated using soft voting.
7 Results and Discussion
The ensemble machine learning models perform better than the individual models. The performance of the ensembles is also superior to the FastText and fine-tuned MuRIL models. The best five performing ensembles along with the FastText and fine-tuned MuRIL models are presented in Table 5 that represent the macro F1 score for each model. Although the scores are encouraging, the performance drops in languages that share a script such as Devanagari for Hindi, Maithili, Marathi, Konkani, and Sanskrit, Arabic for Kashmiri, Urdu, and Sindhi. We compare our models with the state-of-the-art IndicLID model Madhani et al. (2023a) on the Bhasha-Abhijnaanam Madhani et al. (2023a) benchmark of 88K sentences where our model outperforms the IndicLID model Madhani et al. (2023a). The F1-scores of the IndicLID model on the Bhasha-Abhijnaanam dataset is 0.98 as reported in the paper. We could not run their fine-tuned IndicBERT model while the IndicLID FTN model has a macro F1-score of 0.88. Our deep learning baselines perform better on the languages where data are scraped, while the performance dips for languages that are created from other external resources.
7.1 Results for Individual Languages
The language-wise performance of the models is shown in Tables 6. We can observe from the table that the performance drops for extremely low resource languages such as Bodo, Dogri, Maithili, and Sindhi. Deep learning models suffer in identifying those languages for which they have scarcity during pre-training. MuRIL Khanuja et al. (2021) outperforms every other model in languages on which the model was pre-trained.
8 Conclusion
In this paper, we present the ILID benchmark, a manually curated dataset for 23 languages, which includes English and 22 Indian languages containing 250K sentences. It is specifically designed for Indian language identification. Along with the dataset, we develop several machine learning an deep learning models that perform well, even with limited training data. This makes them especially suitable for low-resource languages. To check the consistency and reliability of our ML models, we compare them with deep learning models like MuRIL and FastText. The ILID dataset and models together provide important resources for advancing multilingual NLP research in Indian languages. We hope this work encourages more exploration and development in this area, which has been overlooked.
Lang LR DT RF SVM SGD KNN NB Muril Fasttext
Dev Test Dev Test Dev Test Dev Test Dev Test Dev Test Dev Test Dev Test Dev Test
asm 0.98 0.99 0.96 0.96 0.98 0.98 0.98 0.99 0.97 0.98 0.90 0.96 0.97 0.98 0.99 0.99 0.96 0.96
ben 1.00 1.00 0.97 0.98 0.99 0.99 1.00 1.00 0.99 0.99 0.91 0.97 0.99 0.99 1.00 1.00 0.98 0.98
brx 0.96 0.95 0.91 0.89 0.96 0.94 0.96 0.95 0.94 0.95 0.85 0.93 0.96 0.95 0.91 0.91 0.92 0.91
doi 0.92 0.92 0.83 0.82 0.90 0.91 0.91 0.92 0.88 0.88 0.68 0.81 0.90 0.91 0.93 0.92 0.91 0.91
eng 1.00 1.00 0.99 0.99 1.00 1.00 1.00 1.00 1.00 1.00 0.99 1.00 1.00 0.99 1.00 1.00 0.99 0.99
gom 0.97 0.97 0.84 0.83 0.94 0.95 0.96 0.97 0.90 0.91 0.78 0.90 0.94 0.96 0.97 0.97 0.91 0.92
guj 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.95 1.00 1.00 1.00 0.99 0.99
hin 0.99 0.98 0.93 0.92 0.98 0.97 0.99 0.99 0.97 0.97 0.80 0.92 1.00 0.95 0.99 0.99 0.98 0.98
kan 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.98 1.00 1.00 1.00 0.98 0.97
kas 0.99 0.98 0.94 0.94 0.97 0.98 0.99 0.99 0.94 0.95 0.96 0.96 0.85 0.98 0.99 1.00 0.97 0.96
mai 0.90 0.86 0.83 0.78 0.84 0.84 1.00 0.85 0.83 0.82 0.65 0.73 0.84 0.85 0.89 0.88 0.87 0.86
mal 1.00 1.00 0.99 1.00 1.00 1.00 0.86 1.00 1.00 1.00 1.00 1.00 0.95 1.00 1.00 1.00 0.94 0.94
mar 0.99 0.97 1.00 0.84 0.96 0.96 1.00 0.98 0.91 0.91 0.76 0.91 1.00 0.95 0.98 0.98 0.94 0.94
mni_Beng 1.00 1.00 0.99 0.99 1.00 1.00 0.99 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.99
mni_Mtei 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.91 1.00 0.69 0.71 0.99 0.98
npi 0.92 0.91 0.85 0.84 0.91 0.90 1.00 0.91 0.87 0.86 0.75 0.83 1.00 0.91 0.93 0.92 0.88 0.87
ory 1.00 1.00 1.00 1.00 1.00 1.00 0.92 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
pan 1.00 1.00 0.99 0.99 1.00 1.00 1.00 1.00 1.00 1.00 0.99 1.00 0.97 1.00 1.00 1.00 0.96 0.97
san 1.00 1.00 0.95 0.95 1.00 1.00 1.00 1.00 0.96 0.97 0.60 0.96 1.00 0.97 1.00 1.00 0.98 0.98
sat 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.99 0.92 1.00 0.70 0.72 0.99 0.99
snd_arab 1.00 1.00 0.97 0.98 0.99 1.00 1.00 1.00 0.99 1.00 0.99 0.97 0.99 0.98 1.00 1.00 1.00 0.99
snd_deva 0.94 0.93 0.88 0.85 0.92 0.91 0.94 0.92 0.89 0.90 0.80 0.88 0.92 0.92 0.94 0.94 0.92 0.92
tam 1.00 1.00 0.99 0.99 1.00 1.00 1.00 1.00 0.99 0.99 0.90 1.00 0.99 1.00 1.00 1.00 0.93 0.94
tel 1.00 1.00 0.99 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 0.97 0.97
urd 1.00 0.98 0.93 0.93 0.97 0.97 1.00 0.99 0.96 0.96 0.97 0.97 0.98 0.98 1.00 1.00 0.81 0.79
Table 6: F1-Scores on ILID Dev and Test Sets for each classifier on each language. Highest scores in each language in respective datasets across models are marked in bold